OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)?

OpenStudy (agentjamesbond007):

The slope of a perpendicular line is always -(1/m) where m is the slope

Parth (parthkohli):

Just flip the fraction and bring the negative sign on!

OpenStudy (anonymous):

Sorry added the wrong question. Correct question edited.

OpenStudy (agentjamesbond007):

Take the slope of the line that passes both points. 2-2/5-2 slope gives zero. use y-y1=m(x-x1) choose any coordinate given (we'll use 5,2) y-2=0(x-5) y=2 the equation of your line is y=2

OpenStudy (anonymous):

Oh, now I get it. so youre just making an equestion by adding in the numbers with the x and y.

OpenStudy (agentjamesbond007):

Exactly!

OpenStudy (anonymous):

Thank you for the steps and helping.

OpenStudy (agentjamesbond007):

You are welcome!